OpenStudy (anonymous):

Simplify: i^137 + i^1003

OpenStudy (anonymous):

i = sqrt(-1) i^2 = -1 i^3 =-sqrt(-1) i^4 = 1 137 / 4 = 34 + 1/4 So i^137 = sqrt(-1) 1003/4 = 250 + 3/4 So i^1003 = -sqrt(-1) So sqrt(-1) -sqrt(-1) = 0 Answer: 0